40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 26 22:56:01 2011, kiwamu, Update, LSC, c1iscaux2 and c1iscaux restrated 
    Reply  Wed Apr 27 21:19:25 2011, kiwamu, Update, LSC, LSC whitening for PD1-4 Screenshot-1.png
Message ID: 4577     Entry time: Wed Apr 27 21:19:25 2011     In reply to: 4570
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: LSC whitening for PD1-4 

On the back side of 1Y2 rack I found a cable, CAB-1X2-LSC_7, which is supposed to be connected to the whitening filter was disconnected.

I plugged it back and confirmed that the whitening filter is under control of EPICS.

Now all the gain sliders seem to be working because I can change the amplitude of signals with the sliders.

 

(method)

  To check if the gain sliders are working or not, I intentionally disconnected all the inputs to the whitening filter.

Then I brought a gain slider of interest to the maximum. Due to the big gain I was easily able to see noise lying above ADC noise.

Also if the gain slider is 0 dB, which is the minimum value, the spectrum becomes just ADC noise.

In this way I checked all the gain sliders from PD1 to PD4. The picture below is just an example screenshot when I was doing this test.

Note that each filer is designed to have two poles at 150 Hz and two zeros at 15 Hz.

Screenshot-1.png

Quote from #4570

While checking whitening filters on the LSC rack, I found some epics controls for the whitening looked not working.

So I powered two crates off : the top one and the bottom one on 1Y3 rack.

These crates contain c1iscaux and c1iscaux2. Then powered them on. But it didn't solve the issue.

ELOG V3.1.3-